The oxygen molecule, O2, has a total mass of 5.30×10-26 kg and a rotational inertia of 1.94×10-46kg-m2 about an axis perpendicular to the center of the line joining the atoms. Suppose that such a molecule in a gas has a speed of 1.48×102m/s and that its rotational kinetic energy is two-thirds (2/3) of its translational kinetic energy.
Find its angular velocity.
Answer #1

Mass of the oxygen molecule = m = 5.3 x 10-26 kg

Rotational inertia of the oxygen molecule = I = 1.94 x 10-46 kg.m2

Speed of the molecule = V = 1.48 x 102 m/s

Angular velocity of the molecule =

Translational kinetic energy of the molecule = KET = mV2/2

Rotational kinetic energy of the molecule = KER = I2/2

The rotational energy of the molecule is two-thirds of the translation energy of the molecule.

KER = (2/3)KET

= 1.99 x 1012 rad/s

Angular velocity of the molecule = 1.99 x 1012 rad/s